OpenStudy (anonymous):

PLEASE 5x - 4y = 14 x - 2y = -2

OpenStudy (austinl):

We need to get one of the variables in the bottom equation to be the same as the equation on top. Which variable do you want to do first?

OpenStudy (anonymous):

5?

OpenStudy (austinl):

Ok, the top equation stays the same. 5x-4y=14 but the bottom equation is multiplied by -5. -5(x-2y=-2) which is -5x+10y=10 Which gives us 5x-4y=14 -5x+10y=10

OpenStudy (anonymous):

OMG youre perf

OpenStudy (austinl):

The two would cancel, which would give -4y=14 10y=10 But it occurs to me that there is an easier way to do this... :| oops... lol

OpenStudy (anonymous):

i love you

OpenStudy (austinl):

5x - 4y = 14 x - 2y = -2 I would do this, x=2y-2 5(2y-2)-4y=14 10y-10-4y=14 6y-10=14 6y=24 y=4 Then you would take that, and plug it back into an equation to get the answer.

OpenStudy (austinl):

Think you can handle that?

OpenStudy (anonymous):

yes

OpenStudy (austinl):

Coolio!